Rogen Corporation manufactures a single product. The standard cost per unit of product is shown below. Direct materials—1 pound plastic at $6.00 per pound $ 6.00 Direct labor—2.0 hours at $11.90 per hour 23.80 Variable manufacturing overhead 12.00 Fixed manufacturing overhead 20.00 Total standard cost per unit $61.80 The predetermined manufacturing overhead rate is $16.00 per direct labor hour ($32.00 ÷ 2.0). It was computed from a master manufacturing overhead budget based on normal production of 10,400 direct labor hours (5,200 units) for the month. The master budget showed total variable costs of $62,400 ($6.00 per hour) and total fixed overhead costs of $104,000 ($10.00 per hour). Actual costs for October in producing 4,100 units were as follows. Direct materials (4,240 pounds) $ 26,288 Direct labor (8,030 hours) 97,163 Variable overhead 99,710 Fixed overhead 33,390 Total manufacturing costs $256,551 The purchasing department buys the quantities of raw materials that are expected to be used in production each month. Raw materials inventories, therefore, can be ignored.

(a) Compute all of the materials and labor variances.

Total materials variance $enter a dollar amount select an option

Materials price variance $enter a dollar amount select an option

Materials quantity variance $enter a dollar amount select an option

Total labor variance $enter a dollar amount select an option

Labor price variance $enter a dollar amount select an option

Labor quantity variance $enter a dollar amount select an option

(b) Compute the total overhead variance. Total overhead variance $enter the total overhead variance in dollars select an option
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
4240 $6.20 $26,288.00 4240 $6.00 $25,440.00 4100 $6.00 $24,600.00
$848.00 U $840.00 U
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $848.00 U
Direct material quantity variance $840.00 U
Direct material cost variance $1,688.00 U
Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
8030 $12.10 $97,163.00 8030 $11.90 $95,557.00 8200 $11.90 $97,580.00
$1,606.00 U $2,023.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $1,606.00 U
Direct Labor Efficiency variance $2,023.00 F
Direct labor cost variance $417.00 F

Solution b:

Total overhead variance = Overhead applied - Actual overhead = (4100*$32) - ($99,710 + $33,390) = $1,900 U